# Non-toric brane webs, Calabi–Yau 3-folds, and 5d SCFTs

### v1

@@ 6.4. Calabi–Yau 3-folds from scattering diagrams
The initial scattering diagram is not consistent in general. However, if the asymptotic web of 5-branes with 7-branes $W^{\mathrm{asym}}$ is consistent, it follows from [47] that there exists a unique consistent scattering diagram $\mathfrak{D}$ obtained from $\mathfrak{D}_{\mathrm{in}}$ by adding outgoing walls. Moreover, $\mathfrak{D}$ can be algorithmically computed– see [47, §3.4], [51] and [10, 12] in higher dimensions.
@@
with $\beta(0)\notin\cup_{(\mathfrak{d},f_{\mathfrak{d}})}\mathfrak{d}$ and whose image consists of finitely many line segments $L_{1},L_{2},\dots,L_{N}$, such that no $L_{i}$ is contained in a wall of $\mathfrak{d}$ and each $L_{i}$ is compact except $L_{1}$. Further, we require that each $L_{i}$ is contained in a 2-dimensional cone of the fan defined by $\overline{W}$. Moreover, to each $L_{i}$ is assigned a monomial $\mu_{i}=a_{i}t^{k_{i}}x^{m_{i}}y^{n_{i}}$, where $a_{i}\in\mathbb{Z}$, $k_{i}\in\mathbb{Z}_{\geq 0}$, $(m_{i},n_{i})\in\mathbb{Z}^{2}\setminus\{0\}$ with $\beta^{\prime}(t)=-(m_{i},n_{i})$ on $L_{i}$. We require $a_{1}=1$ and $k_{1}=0$, and we refer to $(m_{1},n_{1})$ as the asymptotic direction of the broken line. Moreover, for every $11\leq i\leq n-$ 1, we require that $L_{i}\cap L_{i+1}$ is contained in a wall $\mathfrak{d}=\mathbb{R}_{\geq 0}(p,q)$ of $\mathfrak{D}$, with coprime $(p,q)\in\mathbb{Z}^{2}\setminus\{0\}$, and that $m_{i+1}$ is a monomial in $\mu_{i}f_{\mathfrak{d}}^{|pn_{i}-qm_{i}|}$. We refer to $\beta(0)$ as the end-point of $\beta$, and $\mu_{N}$ at the final monomial of $\beta$.
@@
$\mathcal{A}\longrightarrow\mathbb{C}[x^{\pm},y^{\pm}][\![t]\!]$ (6.8)
@@
This recovers the description given in [22, (4.4)] of the Calabi–Yau 3-fold M-theory dual to a web of 5-branes with 7-branes where all 7-branes have parallel monodromy-invariant directions, that is, when all the white dots are introduced on a single edge of the dual GTP. In [22, (4.4)], all 5-branes are assumed to end on a 7-brane, so that $n=\sum_{i=1}^{n}a_{i}$, and so the factor $t^{w-\sum_{i=1}^{n}a_{i}}$ in Equation (6.9) does not appear explicitly. In [22], the result is obtained from a D6-brane description in a Type IIA string duality frame, which is only possible in this particular situation where all 7-branes have parallel monodromy-invariant directions.
Refer to caption444444444444
Figure 6.4. Web of 5-branes and its dual polygon in Example 6.13.
@@
Let $\overline{W}^{\mathrm{asym}}$ be the asymptotic web of 5-branes given by the rays $\mathbb{R}_{\geq 0}(1,0)$, $\mathbb{R}_{\geq 0}(0,-1)$ and $\mathbb{R}_{\geq 0}(-1,1)$ all endowed with weight $4$. The corresponding polarized toric log Calabi–Yau surface is $(\mathbb{P}^{2},\partial\mathbb{P}^{2},\mathcal{O}(4))$, with momentum polytope as in Figure 6.4. In this toric situation, the initial scattering diagram consisting of these three rays all endowed with the functions $t^{4}$ is already consistent. Choosing a point $p$ in the upper-right 2-dimensional cone of the polyhedral decomposition of $\mathbb{R}^{2}$ defined by $\overline{W}^{\mathrm{asym}}$, the corresponding theta functions are $\vartheta_{1,0}=x$, $\vartheta_{-1,1}=x^{-1}y$, and $\vartheta_{0,-1}=y^{-1}t^{4}$. It follows that $\overline{\mathcal{X}}^{\mathrm{can}}$ is the toric Calabi–Yau 3-fold defined by the equation
@@
Let $\overline{W}^{\mathrm{asym}}$ be as in Example 6.13, and let $W$ be the web of 5-branes with 7-branes obtained by ending two of the four $(1,0)$ 5-branes on a 7-brane on $\mathbb{R}_{>0}(1,0)$. This web is consistent, see Figure 6.5 for a corresponding Symington polygon. The corresponding log Calabi–Yau surface $(Y,D,L)$ consists of the surface $Y$ obtained from $\mathbb{P}^{2}$ by a non-toric blow-up of a point on the toric boundary divisor, the strict transform $D$ of the toric boundary divisor of $\mathbb{P}^{2}$, and the ample line bundle $L=(\pi^{\star}\mathcal{O}(4))(-2E)$, where $\pi:Y\rightarrow\mathbb{P}^{2}$ is the blow-up morphism and $E$ is the exceptional curve. By Equation (6.6), the initial scattering diagram $\mathfrak{D}_{\mathrm{in}}$ consists of the three incoming walls $(\mathbb{R}_{\geq 0}(1,0),t^{4}(1+t^{-2}x))$, $(\mathbb{R}_{\geq 0}(0,-1),t^{4})$, and $(\mathbb{R}_{\geq 0}(-1,1),t^{4})$. The consistent scattering diagram $\mathfrak{D}$ is obtained from $\mathfrak{D}_{\mathrm{in}}$ by adding the outgoing wall $(\mathbb{R}_{\geq 0}(-1,0),1+t^{2}x)$. Choosing a point $p$ as in Figure 6.5, the corresponding theta functions are $\vartheta_{1,0}=x$, $\vartheta_{-1,1}=x^{-1}y$, and $\vartheta_{0,-1}=y^{-1}t^{4}(1+t^{-2}x)$. Hence, $\mathcal{X}^{\mathrm{can}}$ is the Calabi–Yau 3-fold defined by the equation
@@
This recovers the middle equation in [22, Figure 5].
[…]
Figure 6.5. Scattering diagram and Symington polygon in Example 6.14
[…]
Figure 6.6. Scattering diagram and Symington polygon in Example 6.15.
@@
Let $\overline{W}^{\mathrm{asym}}$ be the asymptotic web of 5-branes given by the rays $\mathbb{R}_{\geq 0}(1,0)$, $\mathbb{R}_{\geq 0}(0,-1)$ and $\mathbb{R}_{\geq 0}(-1,1)$ all endowed with weight $5$. Let $W$ be the web of 5-branes with 7-branes obtained by ending two of the five $(1,0)$ 5-branes on a 7-brane on $\mathbb{R}_{>0}(1,0)$, and two of the five $(0,-1)$ 5-branes on a 7-brane on $\mathbb{R}_{>0}(0,-1)$. This web is consistent, see Figure 6.6 for a corresponding Symington polygon. The corresponding log Calabi–Yau surface $(Y,D,L)$ consists of the surface $Y$ obtained from $\mathbb{P}^{2}$ by the non-toric blow-up of two points on the toric boundary divisor, the strict transform $D$ of the toric boundary divisor of $\mathbb{P}^{2}$, and the ample line bundle $L=(\pi^{\star}\mathcal{O}(5))(-2E_{1}-2E_{2})$, where $\pi:Y\rightarrow\mathbb{P}^{2}$ is the blow-up morphism and $E_{1}$, $E_{2}$ are the exceptional curves. By Equation (6.6), the initial scattering diagram $\mathfrak{D}_{\mathrm{in}}$ consists of the three incoming walls $(\mathbb{R}_{\geq 0}(1,0),t^{5}(1+t^{-2}x))$, $(\mathbb{R}_{\geq 0}(0,-1),t^{5}(1+t^{-2}y^{-1}))$, and $(\mathbb{R}_{\geq 0}(-1,1),t^{5})$. By [51, Figure 1.2], the consistent scattering diagram $\mathfrak{D}$ is obtained from $\mathfrak{D}_{\mathrm{in}}$ by adding the outgoing walls $(\mathbb{R}_{\geq 0}(-1,0),1+t^{3}x)$, $(\mathbb{R}_{\geq 0}(0,1),1+t^{3}y^{-1})$ and $(\mathbb{R}_{\geq 0}(-1,1),1+tx^{-1}y)$. Choosing a point $p$ as in Figure 6.6, the corresponding theta functions are $\vartheta_{1,0}=x$, $\vartheta_{-1,1}=x^{-1}y(1+t^{3}y^{-1})$, and $\vartheta_{0,-1}=y^{-1}t^{5}(1+t^{-2}x)$. We deduce that $y^{-1}=t^{-5}\frac{\vartheta_{0,1}}{1+t^{-2}\vartheta_{1,0}}$, and so that $\mathcal{X}^{\mathrm{can}}$ is the Calabi–Yau 3-fold defined by the equation
@@
Let $\overline{W}^{\mathrm{asym}}$ be the asymptotic web of 5-branes given by the rays $\mathbb{R}_{\geq 0}(1,0)$, $\mathbb{R}_{\geq 0}(0,-1)$ and $\mathbb{R}_{\geq 0}(-1,1)$ all endowed with weight $6$. Let $W$ be the web of 5-branes with 7-branes obtained by ending two pairs of $(1,0)$ 5-branes on two 7-branes on $\mathbb{R}_{>0}(1,0)$, two pairs of $(0,-1)$ 5-branes on two 7-branes on $\mathbb{R}_{>0}(0,-1)$, and two pairs of $(-1,1)$ 5-branes on two 7-branes on $\mathbb{R}_{>0}(-1,1)$. This web is consistent, see Figure 6.7 for a corresponding Symington polygon. Denote by $(Y,D,L)$ the corresponding log Calabi–Yau surface with line bundle. The surface $Y$ is obtained from $\mathbb{P}^{2}$ by blowing up non-torically six points, two on each of the three toric lines, and so is a cubic surface. The divisor $D$ is the strict transform of the three toric lines in $\mathbb{P}^{2}$. Moreover, we have $L=(\pi^{\star}\mathcal{O}(6))(-\sum_{i=1}^{6}2E_{i})$, where $\pi:Y\rightarrow\mathbb{P}^{2}$ is the blow-up morphism, and $(E_{i})_{1\leq i\leq 6}$ are the six exceptional curves.
By Equation (6.6), the initial scattering diagram $\mathfrak{D}_{\mathrm{in}}$ consists of the three incoming walls $(\mathbb{R}_{\geq 0}(1,0),t^{6}(1+t^{-2}x)^{2})$, $(\mathbb{R}_{\geq 0}(0,-1),t^{6}(1+t^{-2}y^{-1})^{2})$, and $(\mathbb{R}_{\geq 0}(-1,1),t^{6}(1+t^{-2}x^{-1}y)^{2})$. By [49], see also [25], the consistent scattering diagram $\mathfrak{D}$ is obtained from $\mathfrak{D}_{\mathrm{in}}$ by adding infinitely many rays of every possible rational slope. Moreover, the function attached to each ray can be explicitly determined. For instance, by [49, Proposition 2.4], the function $f_{\mathfrak{d}}$ attached to the added outgoing ray $\mathbb{R}(1,0)$ is
@@
and the other added rays are determined by an $SL(2,\mathbb{Z})$ symmetry described in [49, Theorem 2.5]. Even though the consistent scattering diagram contains infinitely many rays, generators and relations for the algebra of theta functions can still be determined, and it follows from [49, Theorem 0.1]-[25, Eq. (79)] that $\mathcal{X}^{\mathrm{can}}$ is the Calabi–Yau 3-fold defined by the equation
@@
[…]
Figure 6.7. Scattering diagram and Symington polygon in Example 6.16.
@@ 6.5. Disk worldsheet instantons and BPS states
of marked rational curves $f:C\rightarrow Y$ of class $\beta$, intersecting $D$ at a unique point $x$ and with contact order $(kp,kq)$ along $D$, that is, whose strict transform has contact order $k$ with the exceptional divisor in the surface obtained from $Y$ by the corner blow-up corresponding to adding the ray $\mathbb{R}_{\geq 0}(p,q)$ in the fan of the toric surface $\overline{Y}$. Then, the function $f_{\mathfrak{d}}$ associated to the outgoing ray $\mathfrak{d}=\mathbb{R}_{\geq 0}(p,q)$ in $\mathfrak{D}$ is given by
@@
The precise relation between the consistent scattering diagram $\mathfrak{D}$ and the BPS spectrum of the 4d $\mathcal{N}=2$ theory defined by an M5-brane on $T^{2}\subset U$ can be described as follows. By Theorem [23, Theorem 8.5], the function $f_{\mathfrak{d}}$ associated to the outgoing ray $\mathfrak{d}=\mathbb{R}_{\geq 0}(p,q)$ in $\mathfrak{D}$ can be uniquely written as
@@
with $\Omega_{(kp,kq),\beta}\in\mathbb{Z}$. Then, the integers $\Omega_{(kp,kq),\beta}$ are the BPS indices [44, Eq. (1.1)] near infinity of the Coulomb branch of the 4d $\mathcal{N}=2$ theory defined by an M5-brane on $T^{2}\subset U$, where $(kp,kq)\in\mathbb{Z}^{2}$ is the electromagnetic charge with respect to the low-energy $U(1)$ gauge theory, and $\beta$ encodes the flavor charge. The consistency of the scattering diagram is a manifestation of the Kontsevich–Soibelman wall-crossing formula for BPS indices [72], as explained in [11, 24, 26, 27]. In particular, the fact that the consistent scattering diagram can be arbitrarily complicated is related to the inherent complexity of the BPS spectrum of 4d $\mathcal{N}=2$ theories, see [45] for instance.

### v2

@@ 6.4. Calabi–Yau 3-folds from scattering diagrams
The initial scattering diagram is not consistent in general. However, if the asymptotic web of 5-branes with 7-branes $W^{\mathrm{asym}}$ is consistent, it follows from [50] that there exists a unique consistent scattering diagram $\mathfrak{D}$ obtained from $\mathfrak{D}_{\mathrm{in}}$ by adding outgoing walls. Moreover, $\mathfrak{D}$ can be algorithmically computed– see [50, §3.4], [54] and [13, 15] in higher dimensions.
@@
with $\beta(0)\notin\cup_{(\mathfrak{d},f_{\mathfrak{d}})}\mathfrak{d}$ and whose image consists of finitely many line segments $L_{1},L_{2},\dots,L_{N}$, such that no $L_{i}$ is contained in a wall of $\mathfrak{d}$ and each $L_{i}$ is compact except $L_{1}$. Further, we require that each $L_{i}$ is contained in a 2-dimensional cone of the fan defined by $\overline{W}$. Moreover, to each $L_{i}$ is assigned a monomial $\mu_{i}=a_{i}t^{k_{i}}x^{m_{i}}y^{n_{i}}$, where $a_{i}\in\mathbb{Z}$, $k_{i}\in\mathbb{Z}_{\geq 0}$, $(m_{i},n_{i})\in\mathbb{Z}^{2}\setminus\{0\}$ with $\beta^{\prime}(t)=-(m_{i},n_{i})$ on $L_{i}$. We require $a_{1}=1$ and $k_{1}=0$, and we refer to $(m_{1},n_{1})$ as the asymptotic direction of the broken line. Moreover, for every $1\leq i\leq n-1$, we require that $L_{i}\cap L_{i+1}$ is contained in a wall $\mathfrak{d}=\mathbb{R}_{\geq 0}(p,q)$ of $\mathfrak{D}$, with coprime $(p,q)\in\mathbb{Z}^{2}\setminus\{0\}$, and that $m_{i+1}$ is a monomial in $\mu_{i}f_{\mathfrak{d}}^{|pn_{i}-qm_{i}|}$. We refer to $\beta(0)$ as the end-point of $\beta$, and $\mu_{N}$ at the final monomial of $\beta$.
@@
$\mathcal{A}\lhook\joinrel\longrightarrow\mathbb{C}[x^{\pm},y^{\pm}][\![t]\!]$ (6.8)
@@
This recovers the description given in [25, (4.4)] of the Calabi–Yau 3-fold M-theory dual to a web of 5-branes with 7-branes where all 7-branes have parallel monodromy-invariant directions, that is, when all the white dots are introduced on a single edge of the dual GTP. In [25, (4.4)], all 5-branes are assumed to end on a 7-brane, so that $n=\sum_{i=1}^{n}a_{i}$, and so the factor $t^{w-\sum_{i=1}^{n}a_{i}}$ in Equation (6.9) does not appear explicitly. In [25], the result is obtained from a D6-brane description in a Type IIA string duality frame, which is only possible in this particular situation where all 7-branes have parallel monodromy-invariant directions.
Figure 24. Web of 5-branes and its dual polygon in Example 6.13.
@@
Let $\overline{W}^{\mathrm{asym}}$ be the asymptotic web of 5-branes given by the rays $\mathbb{R}_{\geq 0}(1,0)$, $\mathbb{R}_{\geq 0}(0,-1)$ and $\mathbb{R}_{\geq 0}(-1,1)$ all endowed with weight $4$. The corresponding polarized toric log Calabi–Yau surface is $(\mathbb{P}^{2},\partial\mathbb{P}^{2},\mathcal{O}(4))$, with momentum polytope as in Figure 24. In this toric situation, the initial scattering diagram consisting of these three rays all endowed with the functions $t^{4}$ is already consistent. Choosing a point $p$ in the upper-right 2-dimensional cone of the polyhedral decomposition of $\mathbb{R}^{2}$ defined by $\overline{W}^{\mathrm{asym}}$, the corresponding theta functions are $\vartheta_{1,0}=x$, $\vartheta_{-1,1}=x^{-1}y$, and $\vartheta_{0,-1}=y^{-1}t^{4}$. It follows that $\overline{\mathcal{X}}^{\mathrm{can}}$ is the toric Calabi–Yau 3-fold defined by the equation
@@
Let $\overline{W}^{\mathrm{asym}}$ be as in Example 6.13, and let $W$ be the web of 5-branes with 7-branes obtained by ending two of the four $(1,0)$ 5-branes on a 7-brane on $\mathbb{R}_{>0}(1,0)$. This web is consistent, see Figure 25 for a corresponding Symington polygon. The corresponding log Calabi–Yau surface $(Y,D,L)$ consists of the surface $Y$ obtained from $\mathbb{P}^{2}$ by a non-toric blow-up of a point on the toric boundary divisor, the strict transform $D$ of the toric boundary divisor of $\mathbb{P}^{2}$, and the ample line bundle $L=(\pi^{\star}\mathcal{O}(4))(-2E)$, where $\pi:Y\rightarrow\mathbb{P}^{2}$ is the blow-up morphism and $E$ is the exceptional curve. By Equation (6.6), the initial scattering diagram $\mathfrak{D}_{\mathrm{in}}$ consists of the three incoming walls $(\mathbb{R}_{\geq 0}(1,0),t^{4}(1+t^{-2}x))$, $(\mathbb{R}_{\geq 0}(0,-1),t^{4})$, and $(\mathbb{R}_{\geq 0}(-1,1),t^{4})$. The consistent scattering diagram $\mathfrak{D}$ is obtained from $\mathfrak{D}_{\mathrm{in}}$ by adding the outgoing wall $(\mathbb{R}_{\geq 0}(-1,0),1+t^{2}x)$. Choosing a point $p$ as in Figure 25, the corresponding theta functions are $\vartheta_{1,0}=x$, $\vartheta_{-1,1}=x^{-1}y$, and $\vartheta_{0,-1}=y^{-1}t^{4}(1+t^{-2}x)$. Hence, $\mathcal{X}^{\mathrm{can}}$ is the Calabi–Yau 3-fold defined by the equation
@@
This recovers the middle equation in [25, Figure 5].
Figure 25. Scattering diagram and Symington polygon in Example 6.14
Figure 26. Scattering diagram and Symington polygon in Example 6.15.
@@
Let $\overline{W}^{\mathrm{asym}}$ be the asymptotic web of 5-branes given by the rays $\mathbb{R}_{\geq 0}(1,0)$, $\mathbb{R}_{\geq 0}(0,-1)$ and $\mathbb{R}_{\geq 0}(-1,1)$ all endowed with weight $5$. Let $W$ be the web of 5-branes with 7-branes obtained by ending two of the five $(1,0)$ 5-branes on a 7-brane on $\mathbb{R}_{>0}(1,0)$, and two of the five (0,1)(0,-1) 5-branes on a 7-brane on $\mathbb{R}_{>0}(0,-1)$. This web is consistent, see Figure 26 for a corresponding Symington polygon. The corresponding log Calabi–Yau surface $(Y,D,L)$ consists of the surface $Y$ obtained from $\mathbb{P}^{2}$ by the non-toric blow-up of two points on the toric boundary divisor, the strict transform $D$ of the toric boundary divisor of $\mathbb{P}^{2}$, and the ample line bundle $L=(\pi^{\star}\mathcal{O}(5))(-2E_{1}-2E_{2})$, where $\pi:Y\rightarrow\mathbb{P}^{2}$ is the blow-up morphism and $E_{1}$, $E_{2}$ are the exceptional curves.
By Equation (6.6), the initial scattering diagram $\mathfrak{D}_{\mathrm{in}}$ consists of the three incoming walls $(\mathbb{R}_{\geq 0}(1,0),t^{5}(1+t^{-2}x))$, $(\mathbb{R}_{\geq 0}(0,-1),t^{5}(1+t^{-2}y^{-1}))$, and $(\mathbb{R}_{\geq 0}(-1,1),t^{5})$. By [54, Figure 1.2], the consistent scattering diagram $\mathfrak{D}$ is obtained from $\mathfrak{D}_{\mathrm{in}}$ by adding the outgoing walls $(\mathbb{R}_{\geq 0}(-1,0),1+t^{3}x)$, $(\mathbb{R}_{\geq 0}(0,1),1+t^{3}y^{-1})$ and $(\mathbb{R}_{\geq 0}(-1,1),1+tx^{-1}y)$. Choosing a point $p$ as in Figure 26, the corresponding theta functions are $\vartheta_{1,0}=x$, $\vartheta_{-1,1}=x^{-1}y(1+t^{3}y^{-1})$, and $\vartheta_{0,-1}=y^{-1}t^{5}(1+t^{-2}x)$. We deduce that $y^{-1}=t^{-5}\frac{\vartheta_{0,1}}{1+t^{-2}\vartheta_{1,0}}$, and so that $\mathcal{X}^{\mathrm{can}}$ is the Calabi–Yau 3-fold defined by the equation
@@
Let $\overline{W}^{\mathrm{asym}}$ be the asymptotic web of 5-branes given by the rays $\mathbb{R}_{\geq 0}(1,0)$, $\mathbb{R}_{\geq 0}(0,-1)$ and $\mathbb{R}_{\geq 0}(-1,1)$ all endowed with weight $6$. Let $W$ be the web of 5-branes with 7-branes obtained by ending two pairs of $(1,0)$ 5-branes on two 7-branes on $\mathbb{R}_{>0}(1,0)$, two pairs of (0,1)(0,-1) 5-branes on two 7-branes on $\mathbb{R}_{>0}(0,-1)$, and two pairs of (1,1)(-1,1) 5-branes on two 7-branes on $\mathbb{R}_{>0}(-1,1)$. This web is consistent, see Figure 27 for a corresponding Symington polygon. Denote by $(Y,D,L)$ the corresponding log Calabi–Yau surface with line bundle. The surface $Y$ is obtained from $\mathbb{P}^{2}$ by blowing up non-torically six points, two on each of the three toric lines, and so is a cubic surface. The divisor $D$ is the strict transform of the three toric lines in $\mathbb{P}^{2}$. Moreover, we have $L=(\pi^{\star}\mathcal{O}(6))(-\sum_{i=1}^{6}2E_{i})$, where $\pi:Y\rightarrow\mathbb{P}^{2}$ is the blow-up morphism, and $(E_{i})_{1\leq i\leq 6}$ are the six exceptional curves.
By Equation (6.6), the initial scattering diagram $\mathfrak{D}_{\mathrm{in}}$ consists of the three incoming walls $(\mathbb{R}_{\geq 0}(1,0),t^{6}(1+t^{-2}x)^{2})$, $(\mathbb{R}_{\geq 0}(0,-1),t^{6}(1+t^{-2}y^{-1})^{2})$, and $(\mathbb{R}_{\geq 0}(-1,1),t^{6}(1+t^{-2}x^{-1}y)^{2})$. By [52], see also [28], the consistent scattering diagram $\mathfrak{D}$ is obtained from $\mathfrak{D}_{\mathrm{in}}$ by adding infinitely many rays of every possible rational slope. Moreover, the function attached to each ray can be explicitly determined. For instance, by [52, Proposition 2.4], the function $f_{\mathfrak{d}}$ attached to the added outgoing ray $\mathbb{R}(1,0)$ is
@@
and the other added rays are determined by an $SL(2,\mathbb{Z})$ symmetry described in [52, Theorem 2.5]. Even though the consistent scattering diagram contains infinitely many rays, generators and relations for the algebra of theta functions can still be determined, and it follows from [52, Theorem 0.1]-[28, Eq. (79)] that $\mathcal{X}^{\mathrm{can}}$ is the Calabi–Yau 3-fold defined by the equation
@@
Figure 27. Scattering diagram and Symington polygon in Example 6.16.
@@ 6.5. Disk worldsheet instantons and BPS states
of marked rational curves $f:C\rightarrow Y$ of class $\beta$, intersecting $D$ at a unique point and with contact order $(kp,kq)$ along $D$, that is, whose strict transform has contact order $k$ with the exceptional divisor in the surface obtained from $Y$ by the corner blow-up corresponding to adding the ray $\mathbb{R}_{\geq 0}(p,q)$ in the fan of the toric surface $\overline{Y}$. Then, the function $f_{\mathfrak{d}}$ associated to the outgoing ray $\mathfrak{d}=\mathbb{R}_{\geq 0}(p,q)$ in $\mathfrak{D}$ is given by
@@
The precise relation between the consistent scattering diagram $\mathfrak{D}$ and the BPS spectrum of the 4d $\mathcal{N}=2$ theory defined by an M5-brane on $T^{2}\subset U$ can be described as follows. By Theorem [26, Theorem 8.5], the function $f_{\mathfrak{d}}$ associated to the outgoing ray $\mathfrak{d}=\mathbb{R}_{\geq 0}(p,q)$ in $\mathfrak{D}$ can be uniquely written as
@@
with $\Omega_{(kp,kq),\beta}\in\mathbb{Z}$. Then, the integers $\Omega_{(kp,kq),\beta}$ are the BPS indices [47, Eq. (1.1)] near infinity of the Coulomb branch of the 4d $\mathcal{N}=2$ theory defined by an M5-brane on $T^{2}\subset U$, where $(kp,kq)\in\mathbb{Z}^{2}$ is the electromagnetic charge with respect to the low-energy $U(1)$ gauge theory, and $\beta$ encodes the flavor charge. The consistency of the scattering diagram is a manifestation of the Kontsevich–Soibelman wall-crossing formula for BPS indices [75], as explained in [14, 27, 29, 30]. In particular, the fact that the consistent scattering diagram can be arbitrarily complicated is related to the inherent complexity of the BPS spectrum of 4d $\mathcal{N}=2$ theories, see [48] for instance.